# Transfer learning of state-based potential games for process optimization in decentralized manufacturing systems

### v1

@@ Abstract
This paper presents a novel transfer learning approach in state-based potential games (TL-SbPGs) for enhancing distributed self-optimization in manufacturing systems. The approach focuses on the practical relevant industrial setting where sharing and transferring gained knowledge among similar-behaved players improves the self-learning mechanism in large-scale systems. With TL-SbPGs, the gained knowledge can be reused by other players to optimize their policies, thereby improving the learning outcomes of the players and accelerating the learning process. To accomplish this goal, we develop transfer learning concepts and similarity criteria for players, which offer two distinct settings: (a) predefined similarities between players and (b) dynamically inferred similarities between players during training. We formally prove the applicability of the SbPG framework in transfer learning. Additionally, we introduce an efficient method to determine the optimal timing and weighting of the transfer learning procedure during the training phase. Through experiments on a laboratory-scale testbed, we demonstrate that TL-SbPGs significantly boost production efficiency while reducing power consumption of the production schedules while also outperforming native SbPGs.
@@
Acknowledgement(s)
We wish to express our sincere appreciation to our colleagues from the Department of Automation Technology and Learning Systems at South Westphalia University of Applied Sciences for their valuable insights and feedback during this research. Furthermore, we also express our gratitude to all the contributors of MLPro [^1] https://github.com/fhswf/MLPro/graphs/contributors and MLPro-MPPS [^2] https://github.com/fhswf/MLPro-MPPS/graphs/contributors for generously providing their frameworks.
Declaration of Competing Interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
@@ 1 Introduction
Modern manufacturing systems [1] are increasingly expected to be intelligent, agile, highly flexible, and adaptable, given the rapid and continuous changes in their production processes. To address these demands, game theory (GT) applied in dynamic games for engineering applications [2] emerges as a promising approach. GT’s collaborative characteristics are particularly suitable for self-learning in multi-agent systems (MAS), which require fast adaptation and reconfiguration while each agent’s decision-making affects each other. Within industrial settings, GT-based approaches have demonstrated effectiveness in managing decision-making tasks within MAS, where multiple agents collaborate in a shared environment to tackle multi-objective optimization challenges [3]. Despite its potential advantages, the practical implementation of GT in real manufacturing settings remains relatively limited, primarily due to the vast amounts of required training data and the lengthy duration of training [3]. Particularly, learning from scratch implies being unreasonable without access to highly precise digital system representations. Furthermore, the large-scale nature of most production systems worsens the time required for training [4, 5]. However, there are similarities within modules of the system, such as common objectives, hardware actuation, and control parameters, which can be utilised to reduce training time.
Therefore, in this paper, we propose a novel approach that combines successful GT-based learning with transfer learning methodology by exploiting the similarities within system modules. Our objective is to accelerate training duration and improve production efficiency beyond developed GT-based learning methods. Unlike previous GT-based learning methods, our transfer learning method employs communication signals as in [6] to facilitate knowledge transfer among players with common interests during training. The proposed transfer learning aims to preserve the knowledge gained from solving specific cases and efficiently reuse it in different scenarios, thus saving resources. We claim that transfer learning accelerates player learning by extensively learning from the experiences of other players, as illustrated in Fig. 1.
@@
Research on the implementation of transfer learning in industrial control and GT domains is limited. To our knowledge, transfer learning has yet to be applied in GT, including potential games. However, successful practical applications of transfer learning techniques in dynamic programming offer an understanding of how GT can be enhanced. Examples include unsupervised and supervised deep transfer learning for fault diagnosis [7, 8], transfer learning on evolutionary algorithms [9], and transfer learning in reinforcement learning domains [10]. In these cases, transfer setups typically rely on human decisions considering task similarities and are not straightforwardly executable. Moreover, most applications primarily focus on single-agent scenarios. In contrast, our approach targets multi-agent systems while maintaining a distributed player setup. Furthermore, our transfer learning occurs between training players in a parallel manner, rather than involving external or pre-trained agents. In addition, there is currently a lack of large-scale studies in manufacturing automation about transfer learning in self-optimizing production control systems, which would provide valuable real-world examples.
Our proposed approach contains two primary scenarios. Firstly, we propose a transfer learning framework with the assumption that prior analysis has identified similarities in players’ behaviours. By establishing these similarities at the outset of the learning process, knowledge transfer among players can be pre-determined. Secondly, we consider a scenario where initial information about the similarities between players’ behaviours is unavailable. To address this, we develop a methodology through a radial basis function network for measuring similarities among players and facilitating knowledge transfer between players in the absence of such information. The primary benefit of our approach is in its ability to accelerate the learning of players’ behaviours and performance. This is achieved through knowledge sharing among players, which enables faster convergence towards optimal solutions.
@@
    We propose novel transfer learning methodologies for distributed GT-based optimization, specifically in State-based Potential Games (SbPGs), which is called transfer learning of SbPGs (TL-SbPGs). TL-SbPGs encourage players to exchange their experiences and reuse them to optimize their policies.
@@
    We present two transfer learning approaches with an assumption that the similarities between players are known beforehand, such as (1) the momentum-based (MOM) approach and (2) the sliding window (SW) approach. We provide formal proof demonstrating that the proposed transfer learning approaches meet the requirements of SbPGs.
@@
    We propose a transfer learning approach that addresses scenarios where player similarities are unknown, namely radial basic function networks. This approach can also be independently used for similarity measures.
@@
    We propose a method to determine the timing and weighting of transfer learning between players, as we found these factors to impact the effectiveness of the learning process in TL-SbPGs significantly.
@@
    We apply the proposed transfer learning approaches to a laboratory testbed allowing for distributed self-optimization of bulk good operations. We then evaluate the framework in a larger-scale production system, where we thoroughly compare the proposed approaches of TL-SbPGs (without and with similarity measures) on various production scenarios with immensely encouraging results.
@@ 2.1 Self-learning in production systems
Automation technology has gained significant attention in modern manufacturing systems, especially in the era of digitization, where automated production optimization plays a vital role [1]. Self-learning production systems are a prime example of this application, where systems can autonomously learn from their experiences, make independent decisions, and self-diagnose their conditions [13, 14]. Various methods and applications have been established in this domain, which showcases the effectiveness of automation technology. For instance, Bayesian optimization has been successfully employed for auto-tuning throttle valves [15], knowledge-assisted reinforcement learning has proven beneficial for turbine control in wind farms [16], automated Programmable Logic Controller (PLC) code generation via evolutionary algorithms [17], vision-based reinforcement learning for bucket elevator process optimization [18], and further applications in production systems [14, 19]. Recent research [6, 20, 21] focus on developing self-optimizing production systems using game theory in dynamic games. Building upon those previous research, where communication signals between players are enabled [6], they have yet to fully explore the potential benefits of such communication among actuators in production systems. In this study, we aim to utilise this capability to accelerate the learning process while enhancing performance. Our prior studies provide a robust foundation for the current research.
@@ 2.2 Transfer learning for automation in manufacturing
Transfer learning is an advanced technique in the machine learning area and has emerged as a promising paradigm for enhancing automation in manufacturing systems by leveraging knowledge from related tasks or domains [22, 23]. Transfer learning facilitates the application of knowledge acquired from source tasks to accelerate the learning process of new tasks [10, 24]. In practical scenarios, the concept of transfer learning extends beyond a singular activity, which enables solutions to diverse cases such as cross-domains, cross-environments, or cross-phases [25]. In recent literature, researchers have explored various transfer learning techniques to address the challenges of adapting models trained on source domains to target domains with limited or no labelled data [26]. For instance, in the context of robotics and automation, transfer learning has been applied to tasks such as object recognition [27], manipulation and scheduling [28, 29], as well as motion planning [30], where pre-trained models are fine-tuned or adapted to specific manufacturing environments. Moreover, transfer learning has been utilized to transfer knowledge between different manufacturing processes or production lines, which enables more efficient utilization of resources and faster deployment of automation solutions. However, there are no large-scale studies about transfer learning in self-optimizing production control systems that can represent a real-world example, as demonstrated in our study. Additionally, the majority of applications primarily focus on single-agent scenarios.
@@ 2.3 Potential games
Potential Game (PG) was originally introduced in economics [31], but nowadays has been increasingly applied in engineering tasks [2]. For instance, distributed parameter estimation of sensors in a non-convex problem [32] and maximizing system throughput of transmission rate [33] have been reported. As in [31], the potential function $\phi$ aids in achieving local or global optimality of the learned system and facilitates convergence studies. Additionally, incorporating various learning algorithms can ensure the convergence of a system to a Nash equilibrium (NE). However, the major challenge in PG-based optimization for a self-learning system is establishing a strategy that guarantees an untroubled alignment between the potential function and the local objective functions (utilities $U_{i}$) for each player $i$. Hence, to incorporate state information in dynamic PGs (DPGs) [11], SbPGs have been developed in [34] for the first time. Then, SbPGs have been implemented in distributed production systems as an algorithm for self-optimizing the systems [3]. However, the fact is the learning of the distributed policies is mostly conducted from scratch resulting in lengthy training periods and is barely applicable in the real environment. In [6], they extensively discussed communication and memory within SbPGs, yet the exploration of the full potential of communication remains unexplored. Furthermore, in the related previous work [20], they investigated transfer learning between PLC code and GT players. In contrast, this study diverts its focus to transfer learning exclusively between GT players, which facilitates information exchange during the learning process and incorporates similarity measures. Additionally, an algorithm that integrates online model-based learning with SbPGs has been developed in [21, 35]. These referenced works highlight the significance of SbPGs in manufacturing control.
@@ 3 Problem statement
Figure 2: Considered distributed system structure including several subsystems with their own control and communication system.
Each subsystem interacts with its control system by exchanging local signals. Additionally, there are two connections linking different subsystems, which are one at the communication level and another at the process level. The communication level is facilitated by communication interfaces, such as Ethernet, fieldbus, or wireless connections. These communication interfaces enable modules to be uniquely identified within the production environment, which facilitates knowledge transfer via communication signals in our study. We assume that the transfer learning process occurs in both serial and serial-parallel process chains, as depicted in Fig. 3. Similar to [3], we consider that the production chain is represented by an alternating arrangement of actuators that operate on the process (e.g., motors, pumps, conveyors, pumps) and physical states representing the process status. These actuators are expected to exhibit both discrete and continuous operational behaviours, which form a hybrid actuation system.
@@
We represent and analyze the considered distributed system via graph theory [37], as further explained in [3]. We define a set of actuators $\mathcal{N}={1;\ldots;N}$ and a set of states $\mathcal{S}\subset\mathbb{R}^{m}$ with a corresponding set of actions $\mathcal{A}_{i}\subset\mathbb{R}^{c}\times\mathbb{N}^{d}$ with continuous and/or discrete behaviour. We describe the production chain in the form of a directed graph with $|\mathcal{S}|$ states and $N$ actuator nodes. We assume a dynamic series of states and actuators, such that the set of edges $\mathcal{E}$ omits edges $e=(s_{i},s_{j})$ and $e=(A_{i},A_{j})$ with $s_{i},s_{j}\in\mathcal{S}$ and $A_{i},A_{j}\in\mathcal{N}$. Moreover, two sets of surrounding states are determined for every actuator, including the group of subsequent states $\mathcal{S}^{A_{i}}_{next}=\{s_{j}\in\mathcal{S}|\exists e=(A_{i},s_{j})\in\mathcal{E}\}$ and prior states $\mathcal{S}^{A_{i}}_{prior}=\{s_{j}\in\mathcal{S}|\exists e=(s_{j},A_{i})\in\mathcal{E}\}$. The functions of the local utility is characterized as $U_{i}=f(S^{A_{i}},a_{i})$, in which $a_{i}\in\mathcal{A}_{i}$ and ${S}^{A_{i}}\in\mathcal{S}^{A_{i}}=\mathcal{S}^{A_{i}}_{prior}\cup\mathcal{S}^{A_{i}}_{next}\cup\mathcal{S}^{g}$. The group of states $\mathcal{S}^{g}$ are utilized to determine global objectives.
In our study, we enable communication interfaces to facilitate transfer learning between players via communication signals, which leads to an introduction of an auxiliary utility function $H_{i,j}$ that encourages transfer between player $i$ and player $j$. We intend to subsequently integrate this into the local utility function of each player, a topic to be addressed in Section 5. However, in this setup, the multiple utility functions may share states, and the overarching objective is to maximize overall utility, as outlined below:
@@
by jointly accelerating local utilities $U_{i}(a_{i},S^{A_{i}})$ of the system [11], where $\phi$ denotes a global objective function, known as the potential function. As a general framework of this setup, we utilize SbPGs for distribution optimization, as explained in the next section.
This generalized production scenario finds applications across various fields in the process industry sector, including pharmaceuticals, food production, and chemical plants. Decentralized production lines are prevalent in modern industrial environments. To be noted, the assumption that the production chain comprises only sequences of states and actions is not restrictive. A sequence of more than two states can be reformulated into a common state vector, and similar arguments apply to actions. In the manufacturing context, local utilities can represent multi-strategic goals such as minimizing throughput time, idle time, storage requirements, setup and changeover time of machines, reducing production rejects and reworking, and optimizing power consumption. Practical applications often involve a combination of different multi-objectives, which leads to multi-objective optimization problems.

### v3

@@ Abstract
This paper presents a novel online transfer learning approach in state-based potential games (TL-SbPGs) for distributed self-optimization in manufacturing systems. The approach targets practical industrial scenarios where knowledge sharing among similar players enhances learning in large-scale and decentralized environments. TL-SbPGs enable players to reuse learned policies from others, which improves learning outcomes and accelerates convergence. To accomplish this goal, we develop transfer learning concepts and similarity criteria for players, which offer two distinct settings: (a) predefined similarities between players and (b) dynamically inferred similarities between players during training. The applicability of the SbPG framework to transfer learning is formally established. Furthermore, we present a method to optimize the timing and weighting of knowledge transfer. Experimental results from a laboratory-scale testbed show that TL-SbPGs improve production efficiency and reduce power consumption compared to vanilla SbPGs.
@@ 1 Introduction
Modern manufacturing systems [1] are increasingly expected to be intelligent, agile, highly flexible, and adaptable, given the rapid and continuous changes in their production processes. To address these demands, game theory (GT) applied in dynamic games for engineering applications [2] emerges as a promising approach. GT’s collaborative characteristics are particularly suitable for self-learning in multi-agent systems (MAS), which require fast adaptation and reconfiguration while each agent’s decision-making affects each other. Within industrial settings, GT-based approaches have demonstrated effectiveness in managing decision-making tasks within MAS, where multiple agents collaborate in a shared environment to tackle multi-objective optimization challenges [3]. Despite its potential advantages, the practical implementation of GT in real manufacturing settings remains relatively limited, primarily due to the vast amounts of required training data and the lengthy duration of training [3]. Particularly, learning from scratch is often impractical without access to highly precise digital system representations. Furthermore, the large-scale nature of most production systems worsens the time required for training [4, 5]. However, there are similarities within modules of the system, such as common objectives, hardware actuation, and control parameters, which can be utilized to reduce training time.
Therefore, in this paper, we propose the first approach that applies online transfer learning within state-based potential games (SbPGs) for distributed manufacturing systems, combining established GT learning with a similarity-based knowledge transfer mechanism across system modules. Our objective is to accelerate training duration and improve production efficiency beyond developed GT-based learning methods. Unlike previous GT-based learning methods, our transfer learning method employs communication signals as in [6] to facilitate knowledge transfer among players with common interests during training. Our approach aims to preserve the knowledge gained from solving specific cases and efficiently reuse it in different scenarios, thus saving resources. We claim that transfer learning accelerates player learning by extensively learning from the experiences of other players, as illustrated in Fig. 1.
@@
Research on the implementation of transfer learning in industrial control and GT domains is limited. To our knowledge, transfer learning has not yet been applied in GT, including potential games. However, successful practical applications of transfer learning techniques in dynamic programming offer an understanding of how GT can be enhanced. Examples include unsupervised and supervised deep transfer learning for fault diagnosis [7, 8], transfer learning on evolutionary algorithms [9], and transfer learning in reinforcement learning domains [10]. In these cases, transfer setups typically rely on human decisions considering task similarities and are not straightforwardly executable. Moreover, most applications primarily focus on single-agent scenarios. In contrast, our approach targets multi-agent systems while maintaining a distributed player setup. Furthermore, our transfer learning occurs between training players in a parallel manner, rather than involving external or pre-trained agents. In addition, there is currently a lack of large-scale studies in manufacturing automation about transfer learning in self-optimizing production control systems, which would provide valuable real-world examples.
Our proposed approach contains two primary scenarios. Firstly, we propose a transfer learning framework with the assumption that prior analysis has identified similarities in players’ behaviors. By establishing these similarities at the outset of the learning process, knowledge transfer among players can be pre-determined. Secondly, we consider a scenario where initial information about the similarities between players’ behaviors is unavailable. To address this, we develop a methodology through a radial basis function network for measuring similarities among players and facilitating knowledge transfer between players in the absence of such information. The primary benefit of our approach is its ability to accelerate the learning of players’ behaviors and performance. This is achieved through knowledge sharing among players, which enables faster convergence towards optimal solutions.
@@
    We introduce TL-SbPGs, a novel transfer learning methodology for distributed optimization in SbPGs, which enables players to exchange and reuse experiences to optimize their policies.
@@
    We present two transfer learning approaches based on predefined player similarities: (1) the momentum-based (MOM) approach and (2) the sliding window (SW) approach. We provide formal proof that both approaches satisfy the requirements of SbPGs.
@@
    We propose an approach to handle unknown player similarities, using radial basis function networks. This method can also be applied independently for similarity measurement.
@@
    We introduce a method to optimize the timing and weighting of transfer learning, which demonstrates that these factors significantly affect the effectiveness of TL-SbPGs.
@@
    We validate the proposed methodologies through experiments on a laboratory testbed and larger-scale production systems, which show that TL-SbPGs with and without similarity measures deliver highly promising results across various production scenarios.
@@ 2.1 Self-learning in production systems
Automation technology has gained significant attention in modern manufacturing systems, especially in the era of digitization, where automated production optimization plays a vital role [1]. Self-learning production systems are a prime example of this application, where systems can autonomously learn from their experiences, make independent decisions, and self-diagnose their conditions [13, 14]. Various methods and applications have been established in this domain, which showcase the effectiveness of automation technology. For instance, Bayesian optimization has been successfully employed for auto-tuning throttle valves [15], knowledge-assisted reinforcement learning has proven beneficial for turbine control in wind farms [16], automated Programmable Logic Controller (PLC) code generation via evolutionary algorithms [17], vision-based reinforcement learning for bucket elevator process optimization [18], and further applications in production systems [14, 19].
Recent research [20, 21, 22] has shown that GT in dynamic games outperforms multi-agent reinforcement learning in self-learning for distributed manufacturing systems, due to its cooperative decision-making strategies. For example, in [23, 24], they introduce deep multi-agent reinforcement learning with a centralized critic and a central training, decentralized deployment paradigm. While this approach requires no communication during operation, it assumes full access to the state space during training, which is an unrealistic condition in real-world applications. Actor-critic frameworks, as in [25], also fall short compared to GT-based learning due to limited coordination among agents and the absence of guaranteed convergence to a Nash equilibrium. In addition, building upon the previous research, where communication signals between players in GT are enabled [6], they have yet to fully explore the potential benefits of such communication among actuators in production systems. In this study, we aim to utilize this capability to accelerate the learning process while enhancing performance. Our prior studies provide a robust foundation for the current research. Despite the promise of self-learning systems, effectively reusing prior knowledge remains challenging, which motivates our exploration of transfer learning in the following subsection.
@@ 2.2 Transfer learning for automation in manufacturing
Transfer learning is an advanced technique in the machine learning area and has emerged as a promising paradigm for enhancing automation in manufacturing systems by leveraging knowledge from related tasks or domains [26, 27]. Transfer learning facilitates the application of knowledge acquired from source tasks to accelerate the learning process of new tasks [10, 28]. This approach has proven effective in domains like medical imaging and cybersecurity, where transfer learning and explainable artificial intelligence support real-world decision systems through feature reuse, domain adaptation, training efficiency, and interpretability [29, 30]. In practical scenarios, the concept of transfer learning extends beyond a singular activity, which enables solutions to diverse cases such as cross-domains, cross-environments, or cross-phases [31]. In recent literature, researchers have explored various transfer learning techniques to address the challenges of adapting models trained on source domains to target domains with limited or no labelled data [32]. For instance, in the context of robotics and automation, transfer learning has been applied to tasks such as object recognition [33], manipulation and scheduling [34, 35], as well as motion planning [36], where pre-trained models are fine-tuned or adapted to specific manufacturing environments.
Another transfer learning approach involves teacher-student learning, where traditional operations research solvers act as teachers to guide reinforcement learning algorithms in solving robot scheduling problems [37]. Moreover, transfer learning has also been utilized to transfer knowledge between different manufacturing processes or production lines, which enables more efficient utilization of resources and faster deployment of automation solutions. However, there are no large-scale studies about transfer learning in self-optimizing production control systems that can represent a real-world example, as demonstrated in our study. Additionally, the majority of applications primarily focus on single-agent and offline scenarios. In contrast, meta-learning aims to rapid adaptation to new environments by generalizing across tasks [38], but it requires extensive meta-training and assumes consistent task distributions [39], which limits its use in dynamic and heterogeneous manufacturing settings. Our GT-based approach avoids these assumptions and adapts via inter-agent coordination without prior task knowledge.
@@ 2.3 Potential games
Potential Game (PG) was originally introduced in economics [40], but nowadays has been increasingly applied in engineering tasks [2]. For instance, distributed parameter estimation of sensors in a non-convex problem [41] and maximizing system throughput of transmission rate [42] have been reported. As in [40], the potential function $\phi$ aids in achieving local or global optimality of the learned system and facilitates convergence studies. Additionally, incorporating various learning algorithms can ensure the convergence of a system to a Nash equilibrium (NE). However, the major challenge in PG-based optimization for a self-learning system is establishing a strategy that guarantees an untroubled alignment between the potential function and the local objective functions (utilities $U_{i}$) for each player $i$. Hence, to incorporate state information in dynamic PGs (DPGs) [11], SbPGs have been developed in [43] for the first time. Then, SbPGs have been implemented in distributed production systems as an algorithm for self-optimizing the systems [3]. However, the fact is that the learning of the distributed policies is mostly conducted from scratch, resulting in lengthy training periods and is barely applicable in the real environment. In [6], they extensively discussed communication and memory within SbPGs, yet the exploration of the full potential of communication remains unexplored. Additionally, an algorithm that integrates online model-based learning within SbPGs has been developed in [21, 44]. Furthermore, in previous related work [20], transfer learning between PLC code and GT players has been explored using a teacher-student learning framework, where the PLC code acted as a teacher guiding GT players to mimic its behaviour. Existing transfer learning techniques typically assume the presence of a pre-trained agent, whose knowledge is transferred to other agents after its training is completed. In contrast, this study focuses on transfer learning exclusively between GT players, which implements an online approach that enables information exchange during the learning process and incorporates similarity measures.
Given the advantages and challenges of PGs and SbPGs in distributed optimization, we now define the specific problem addressed in this work, centered around modular and decentralized manufacturing systems.
@@ 3 Problem statement
Figure 2: Considered a distributed system structure including several subsystems with their own control and communication systems.
Each subsystem interacts with its control system by exchanging local signals. Additionally, there are two connections linking different subsystems, which are one at the communication level and another at the process level. The communication level is facilitated by communication interfaces, such as Ethernet, fieldbus, or wireless connections. These communication interfaces enable modules to be uniquely identified within the production environment, which facilitates knowledge transfer via communication signals in our study. We assume that the transfer learning process occurs in both serial and serial-parallel process chains, as depicted in Fig. 3. Similar to [3], we consider that the production chain is represented by an alternating arrangement of actuators that operate on the process (e.g., motors, pumps, conveyors, pumps) and physical states representing the process status. These actuators are expected to exhibit both discrete and continuous operational behaviors, which form a hybrid actuation system.
@@
We represent and analyze the considered distributed system via graph theory [46], as further explained in [3]. We define a set of actuators $\mathcal{N}={1;\ldots;N}$ and a set of states $\mathcal{S}\subset\mathbb{R}^{m}$ with a corresponding set of actions $\mathcal{A}_{i}\subset\mathbb{R}^{c}\times\mathbb{N}^{d}$ with continuous and/or discrete behaviour. We describe the production chain in the form of a directed graph with $|\mathcal{S}|$ states and $N$ actuator nodes. We assume a dynamic series of states and actuators, such that the set of edges $\mathcal{E}$ omits edges $e=(s_{i},s_{j})$ and $e=(A_{i},A_{j})$ with $s_{i},s_{j}\in\mathcal{S}$ and $A_{i},A_{j}\in\mathcal{N}$. Moreover, two sets of surrounding states are determined for every actuator, including the group of subsequent states $\mathcal{S}^{A_{i}}_{next}=\{s_{j}\in\mathcal{S}|\exists e=(A_{i},s_{j})\in\mathcal{E}\}$ and prior states $\mathcal{S}^{A_{i}}_{prior}=\{s_{j}\in\mathcal{S}|\exists e=(s_{j},A_{i})\in\mathcal{E}\}$. The functions of the local utility is characterized as $U_{i}=f(S^{A_{i}},a_{i})$, in which $a_{i}\in\mathcal{A}_{i}$ and ${S}^{A_{i}}\in\mathcal{S}^{A_{i}}=\mathcal{S}^{A_{i}}_{prior}\cup\mathcal{S}^{A_{i}}_{next}\cup\mathcal{S}^{g}$. The group of states $\mathcal{S}^{g}$ is utilized to determine global objectives.
In our study, we enable communication interfaces to facilitate transfer learning between players via communication signals, which leads to the introduction of an auxiliary utility function $H_{i,j}$ that encourages transfer between player $i$ and player $j$. We intend to subsequently integrate this into the local utility function of each player, a topic to be addressed in Section 5. However, in this setup, the multiple utility functions may share states, and the overarching objective is to maximize overall utility, as outlined below:
@@
by jointly accelerating local utilities $U_{i}(a_{i},S^{A_{i}})$ of the system [11], where $\phi$ denotes a global objective function, known as the potential function. As a general framework for this setup, we utilize SbPGs for distribution optimization, as explained in the next section.
This generalized production scenario is applicable across industries such as pharmaceuticals, food production, and chemical plants, as well as in energy systems like smart grids and logistics, provided the problem setup involves distributed optimization, as explained in [47]. Decentralized production lines, which require flexible and scalable solutions, are increasingly common in modern industrial environments. To be noted, the assumption that the production chain comprises only sequences of states and actions is not restrictive. A sequence of more than two states can be reformulated into a common state vector, and similar arguments apply to actions. In the manufacturing context, local utilities can represent multi-strategic goals such as minimizing throughput time, idle time, storage requirements, setup and changeover time of machines, reducing production rejects and rework, and optimizing power consumption. Practical applications often involve a combination of different multi-objectives, which leads to multi-objective optimization problems.
To effectively model the distributed control problem formulated above, we apply the SbPGs, whose key concepts and formal properties are detailed in the following section.